OpenStudy (anonymous):

someone help? (x ^2 - 4)÷ (x - 1)

OpenStudy (anonymous):

\[\frac{x^2-4}{x-1}\]?

OpenStudy (anonymous):

I put the answer on your other thread

OpenStudy (anonymous):

it does not divide evenly for sure. you get \[\frac{(x+2)(x-2)}{x-1}\] if you want to divide using synthetic division you will get \[x+1-\frac{3}{x-1}\]

OpenStudy (anonymous):

could you help me with one more please? :(